# Physics-informed renormalisation group flows

### v2

@@ IV.2 Regulator dependences and optimisation in feed-down flows
On the basis of the explicit example in the derivative expansion in the last Section we provide further insight in feed-down effective actions and more generally target actions $\Gamma_{T}\left[\phi[\varphi]\right]$. We first remark that the full effective action $\Gamma_{\varphi}[\varphi]$ is regulator-independent,
$\frac{\delta\Gamma_{\varphi}[\varphi]}{\delta R_{k}}=0\,.$ (94)
Equation 94 is also the cornerstone of RG-consistency, see [6, 36, 43]. Importantly, the property 94 is absent for generalised flows: The infinitesimal transformation $\dot{\phi}[\phi,R]$ carries an explicit $R_{k}$ dependence and so does $\phi_{k=0}[\varphi,R]$. Therefore, the generating functionals $Z_{k=0}[J_{\phi},R]$ and $\Gamma_{T}[\phi[\varphi,R],R]$ depend on the trajectory $R$. We infer from this, that the regulator dependence of expansion coefficients of $\Gamma_{T}[\phi[\varphi]$ is not a good convergence test and systematic error estimate. However, observables are regulator-independent and we can extract convergence properties and error estimates. This is a direct consequence of the reformulation of the task to compute the effective action $\Gamma_{\varphi}$ to that of computing the pair $(\Gamma_{T},\phi)$. In Section III.2 we have already discussed the consequences for functional optimisation.
@@ V Applications of PIRG flows
In Section V.2 we use feed-down flows for the first order of the derivative expansion. This first application showcases the power of PIRG flows and the derived concept of feed-down flows: We can utilise the advances and results obtained so far in a given approximation scheme, and go beyond the respective order without any modification (back-feeding) of the former results. The effective potential is computed with the standard Wetterich flow, resulting in the well-known results in LPA. The field-dependent wave function is absorbed in the order $\partial^{2}$ part of the flowing field such that the anomalous dimension does not feed back into the effective potential. We provide solutions at the critical point and in the broken phase and compare to the ground state expansion 17 or [2].
@@ V.1.2 Classical target action flow in O(1)
$\left[1+\frac{2}{d+2}\dot{\phi}^{\prime}\right]\frac{A_{d}k^{d+2}}{k^{2}+\mu_{\varphi}+\frac{3}{2}\lambda_{\varphi}\phi^{2}}-\dot{\cal C}_{k}\,,$ (106)
@@
[…]
×$\hskip 22.76219pt\times\exp\left(f\left(\frac{\phi^{2}}{2}\right)-f\left(\frac{x^{2}}{2}\right)\right)\,,$ (108a)
@@
The upper integration boundary of 108 corresponds to the boundary condition of the ODE and is set to $\infty$ to implement locality as discussed in Section II.3.1 (a). Then, the transformation $\dot{\phi}$ is always well defined and finite: Due to the integration boundary, the integration variable $y$ is always bigger than $\phi$, which turns the exponential function in the integrand $<1$ for $\phi\geq 0$. Furthermore, we fix the constant $c_{1}$ to implement a $\phi\to-\phi$ symmetry in $\dot{\phi}$. For any other choice $\dot{\phi}$ diverges as $\phi\to-\infty$, this is discussed further in Appendix A.
@@ V.2 First order derivative expansion in the O(1)-model with the $V_{W}$ target action
where $V_{W}$ is the fully $k$ dependent LPA potential. It is computed with a simple finite difference implementation, which was outlined in [50]. Alternatively, we can also use an implementation in terms of Continuous or Discontinuous Galerkin methods, which allow us to capture the intricate dynamics in the potential which is linked to first order phase transitions [45, 52, 49]. Our present developments utilise simple, ready-to-use codes which are tailor-made for the LPA task and beyond. Codes are openly accessible in the GitHub repositories linked to [49, 50].
@@ V.2.2 Results for the O(1)-model
$2f\ V_{W}^{(1)}=$ $\int\frac{d^{d}q}{(2\pi)^{d}}R_{k}(q)G_{W}(q)\times\left[2f^{(1)}-2fV_{W}^{(3)}\frac{\partial}{\partial p^{2}}\Big{\{}\left.G_{W}(q+p)\,(p+q)\cdot p+G_{W}(q-p)\,(p-q)\cdot p\Big{\}}\right|_{p=0}\right]$
$+\int\frac{d^{d}q}{(2\pi)^{d}}\Biggl{\{}\partial_{t}R_{k}(q)G^{2}_{W}(q)\left(V_{W}^{(3)}\right)^{2}\frac{\partial}{\partial p^{2}}\Big{\{}\left.G_{W}(p+q)\Big{\}}\right|_{p=0}\Biggr{\}}\,,$ (124)
@@
and varying initial masses $m_{\varphi}^{2}<0$ for $d=3$. The RG-time integration is performed up to $t=-20$, i.e. $k<10^{-8}$ with an implicit QBDF time-stepping algorithm [55].
@@ VI Conclusion
In this work we have have suggested and worked out a novel functional renormalisation group approach with physics-informed renormalisation group flows (PIRG flows), which is detailed in Section III. The great potential of this novel approach has been illustrated at the example of key applications, and is based on a radical shift of perspective on RG flows: Instead of seeing RG flows as a mere tool for computing the effective action $\Gamma_{\varphi}[\varphi]$ of the fundamental field $\varphi$, they give access to pairs
@@
The functional renormalisation group approach with PIRG flows offers a novel RG approach, whose flexibility allows for the most direct access to physics phenomena and mechanisms as well as a significant minimisation of the involved computational costs. Unlocking the great potential of PIRG flows should allow us to significantly further our understanding and quantitative access to the physics of strongly correlated systems. There are many applications that are natural extensions of that presented here. Here we only name a few of them that are already under investigation: (1) The reparametrisation potential of PIRG flows can be used to construct fully gauge invariant flows, and this possibility originates in the very fact that the composite field effective action is not identical to that in the original fields. (2) Ground state expansions are highly interesting for an application to non-equilibrium evolutions of closed and open quantum systems, including systems with an initial far-from-equilibrium state. One may even base expansion schemes on time-dependent ground state expansions. (3) The flexibility of the choice of the target action allows for the systematic inclusion of results from other approaches such as (resummed) perturbation theory, hard thermal and dense loop results of low and high order, or density functional theory. (4) Finally, a combination of several of the present advances will be applied to the ongoing endeavour of resolving the physics of QCD at finite temperature and density. A specific first application is the phase structure of QCD, including the location of the critical end point with full apparent convergence and a small systematic error. We hope to report on these applications in the near future.
@@ Appendix A Symmetries and cumulants-preserving physics-informed RG flows
In this Appendix we discuss the choice of the constant part $\dot{\mathcal{C}}$ of the PIRG flow, see 66. It carries the $k$ dependence of the normalisation $\mathcal{N}_{\phi}$ of the path integral 6, as well as genuine deformations of the field $\Delta\dot{\mathcal{C}}_{k}$. In Section A.1, we show, how the rather formal cumulants-preserving property 2 turns into a practical cumulants-preserving property for $\dot{\cal C}$, based on the triviality property of the PIRG for asymptotically large composite fields. In Section A.2 the cumulants-preserving property of the flow is used to disentangle the physical part of ${\dot{C}}$ and the normalisation flow $\partial_{t}\ln{\cal N}_{k}$. In Sections A.4 and A.3 we illustrate some of the specifics at the example of the free theory and discuss symmetry properties of the classical target action flows. The latter allows us to further simplify the cumulants-preserving flows. The conceptual findings of this Appendix are then used for the computation of observables in Appendix B, which offers a comprehensive benchmark test.
@@ A.1 Cumulants-preserving PIRGs
$\Gamma_{T}[\phi\to\infty]=\Gamma_{T,\textrm{as}}[\phi]=S_{\textrm{cl}}[\varphi]+\ln{\cal N}_{k}(\boldsymbol{g})\,,$ (131)
@@ A.1.1 Cumulants-preserving Wetterich flow
Evidently, not both normalisations, ${\cal N}_{k},{\cal N}^{\prime}_{k}$ can be $\boldsymbol{g}$-independent and can satisfy the cumulants-preserving property 60. Equation 137 indeed entails that $\log{\cal N}^{\prime}_{k}$ is $\boldsymbol{g}$-dependent, as the implicit normalisation, induced by the Wetterich flow is not: For the respective proof we use that the effective action reduces to the classical one for large amplitudes, $\varphi\to\infty$,
$\Gamma_{\varphi,k}[\varphi\to\infty]=S_{\textrm{\tiny{UV}}}[\varphi]+\ln{\cal N}_{W,k}\,,$ (138)
@@
Moreover, for $k\to\infty$, all theories tend towards the Gaußian theory with the momentum-dependent mass $R_{k}$, and the normalisation there is simply ${\cal N}_{W,k\to\infty}=1$. In combination this entails
@@ A.1.2 Cumulants-preserving PIRG flow
with a $(\boldsymbol{g}$-independent constant on the right-hand side of 146. Finally we remark, that although the coupling dependence in 146 has to cancel, the specific value of the single terms $\dot{\mathcal{C}}_{k}(\boldsymbol{g})$ and $\Delta\dot{\mathcal{C}}_{k}(\boldsymbol{g})$ is not fixed for a specific set of coupling parameters $\boldsymbol{g}$. This freedom opens the gateway for optimisation procedures. The condition used in the present work is discussed in the following Section.
@@ A.4 Symmetry of the field transformation in the classical target action flow
Figure 8: Flowing field transformation $\partial_{t}\phi=\dot{\phi}$ for different choices of boundary conditions. One corresponds to the upper integration bound in 108 which is set to +$\infty$+ ∞ (or $\infty$- ∞ for the dashed lines), whereas $c_{1}$ parametrises the constant part of the flow 107. It implements the choice of constant discussed in Appendix A. The blue curve indicates the value of $c_{1}$ for which $\dot{\phi}(\phi=\pm\infty)=0$ and 148 are fulfilled. The constant is a scale-dependent quantity; the figure uses $k=0.5$ as an example for the initial conditions specified in Section V.1.
@@ B.1 Reference solution
[…]
@@
[…]
@@
$=-\frac{2}{\mu_{\varphi}}\langle\hat{\varphi}^{2}\rangle^{(c)}-\frac{8x}{\mu_{\varphi}}\frac{K_{\frac{5}{4}}\left(x\right)}{K_{\frac{1}{4}}\left(x\right)}\partial_{x}\log\frac{K_{\frac{5}{4}}\left(x\right)}{K_{\frac{1}{4}}\left(x\right)}\,,$ (160)
@@ B.2 Benchmark with PIRG flows
In the present case we consider cumulants-preserving variants of classical target action, recall the discussion in Appendix A. In particular we implement 148, for which we have verified 146. The computation is initiated at $\Lambda=\exp(5)\approx 150$ and the RG-time is integrated up to $t=-15$, corresponding to $k_{\mathrm{fin}}=4\times 10^{-5}$. The map $\phi(\varphi)$ is reconstructed using 55 and depicted in Figure 9(a). Using the map we can also depict the potential in terms of the original field $V_{\mathrm{cl}}(\phi(\varphi))$. This is shown in Figure 9(b), where we have explicitly omitted the constant shift in the target action 96 for better visibility. One can clearly discern 68 in the Figure: The solutions for the effective actions are genuinely different for different composites $\hat{\phi}$.
@@
which is an algebraic equation in $f_{0}$. We immediately find that for the classical target action flow, i.e. constant $\partial_{t}V_{T}$, $f_{0}$ has a non-trivial singularity where $02\rho V_{T}^{\prime}-\frac{2}{d+2}\frac{A_{d}k^{d+2}}{k^{2}+V_{T}^{\prime}}=$ 0. This singularity can be lifted by choosing $\partial_{t}V_{T}$ accordingly
@@ Appendix E Feed-down flows in the first order of the derivative expansion for O(N)-theories
$+\frac{\phi^{a}\phi^{b}}{2\rho}\,\frac{1}{k^{2}+V^{\prime}_{W}(\rho)+2\rho V^{\prime\prime}_{W}(\rho)}\,.$ (182)

### v3

@@ IV.2 Regulator dependences and optimisation in feed-down flows
On the basis of the explicit example in the derivative expansion in the last Section we provide further insight in feed-down effective actions and more generally target actions $\Gamma_{T}\left[\phi[\varphi]\right]$. On the basis of the explicit example in the derivative expansion in the last Section we provide further insight in feed-down effective actions and more generally target actions $\Gamma_{T}\left[\phi[\varphi]\right]$. We first remark that the full effective action $\Gamma_{\varphi}[\varphi]$ at $k=0$ is regulator-independent,
$\frac{\delta\Gamma_{\varphi}[\varphi]}{\delta R_{\phi}}=0\,.$ (94)
Equation 94 is also the cornerstone of RG-consistency, see [6, 41, 48]. Importantly, the property 94 is absent for generalised flows: The infinitesimal transformation $\dot{\phi}[\phi,R]$ carries an explicit $R_{\phi}$ dependence and so does $\phi_{k=0}[\varphi,R]$. Therefore, the generating functionals $Z_{k=0}[J_{\phi},R]$ and $\Gamma_{T}[\phi[\varphi,R],R]$ depend on the trajectory $R$. We infer from this, that the regulator dependence of expansion coefficients of $\Gamma_{T}[\phi[\varphi]]$ is not a good convergence test and systematic error estimate. However, observables are regulator-independent and we can extract convergence properties and error estimates. This is a direct consequence of the reformulation of the task to compute the effective action $\Gamma_{\varphi}$ to that of computing the pair $(\Gamma_{T},\phi)$. In Section III.2 we have already discussed the consequences for functional optimisation.
@@ V Applications of PIRG flows
In Section V.2 we use feed-down flows for the first order of the derivative expansion. This first application showcases the power of PIRG flows and the derived concept of feed-down flows: We can utilise the advances and results obtained so far in a given approximation scheme, and go beyond the respective order without any modification (back-feeding) of the former results. The effective potential is computed with the standard Wetterich flow, resulting in the well-known results in LPA. The field-dependent wave function is absorbed in the order $\partial^{2}$ part of the flowing field such that the anomalous dimension does not feed back into the effective potential. We provide solutions at the critical point and in the broken phase and compare to the ground state expansion with 17 or [2].
@@ V.1.2 Classical target action flow in O(1)
$\hskip-48.36958pt\left[1+\frac{2}{d+2}\dot{\phi}^{\prime}\right]\frac{A_{d}k^{d+2}}{k^{2}+\mu_{\varphi}+\frac{3}{2}\lambda_{\varphi}\phi^{2}}-\dot{\cal C}_{k}\,$, (106)
@@
$\dot{\phi}=$ $\int_{\phi}^{\infty}dx\,\frac{d+2}{2}\left(1-c_{1}\frac{k^{2}+\mu_{\varphi}+3\lambda_{\varphi}\frac{x^{2}}{2}}{k^{2}}\right)$
$\hskip 22.76219pt\times\exp\left(f\left(\frac{\phi^{2}}{2}\right)-f\left(\frac{x^{2}}{2}\right)\right)\,$, (108a)
@@
The upper integration boundary of 108 corresponds to the boundary condition of the ODE and is set to $\infty$ to implement locality as discussed in Section II.3.1 (a). Then, the transformation $\dot{\phi}$ is always well defined and finite: Due to the integration boundary, the integration variable $y$ is always bigger than $\phi$, which turns the exponential function in the integrand <1<1 for $\phi\geq 0$. Furthermore, we fix the constant $c_{1}$ to implement a $\phi\to-\phi$ symmetry in $\dot{\phi}$. For any other choice $\dot{\phi}$ diverges as $\phi\to$-$\infty$, this is discussed further in Appendix A.
@@ V.2 First order derivative expansion in the O(1)-model with the $V_{W}$ target action
where $V_{W}$ is the fully $k$-dependent LPA potential. It is computed with a simple finite difference implementation, which was outlined in [55]. Alternatively, we can also use an implementation in terms of Continuous or Discontinuous Galerkin methods, which allow us to capture the intricate dynamics in the potential which is linked to first order phase transitions [50, 57, 54]. Our present developments utilise simple, ready-to-use codes which are tailor-made for the LPA task and beyond. Codes are openly accessible in the GitHub repositories linked to [54, 55].
@@ V.2.2 Results for the O(1)-model
$2f\ V_{W}^{(1)}=$ $\int\frac{d^{d}q}{(2\pi)^{d}}R_{\phi}(q)G_{W}(q)\times\left[2f^{(1)}-2fV_{W}^{(3)}\frac{\partial}{\partial p^{2}}\Big{\{}\left.G_{W}(q+p)\,(p+q)\cdot p+G_{W}(q-p)\,(p-q)\cdot p\Big{\}}\right|_{p=0}\right]$
$+\int\frac{d^{d}q}{(2\pi)^{d}}\Biggl{\{}\partial_{t}R_{\phi}(q)G^{2}_{W}(q)\left(V_{W}^{(3)}\right)^{2}\frac{\partial}{\partial p^{2}}\Big{\{}\left.G_{W}(p+q)\Big{\}}\right|_{p=0}\Biggr{\}}\,$, (124)
@@
and varying initial masses $m_{\varphi}^{2}<0$ for $d=3$. The RG-time integration is performed up to t=20t=-20, i.e. $k<10^{-8}$ with an implicit QBDF time-stepping algorithm [60].
@@ VI Conclusion
In this work we have suggested and worked out a novel functional renormalisation group approach with physics-informed renormalisation group flows (PIRG flows), which is detailed in Section III. The great potential of this novel approach has been illustrated at the example of key applications, and is based on a radical shift of perspective on RG flows: Instead of seeing RG flows as a mere tool for computing the effective action $\Gamma_{\varphi}[\varphi]$ of the fundamental field $\varphi$, they give access to pairs
@@
The functional renormalisation group approach with PIRG flows offers a novel RG approach, whose flexibility allows for the most direct access to physics phenomena and mechanisms as well as a significant minimisation of the involved computational costs. Unlocking the great potential of PIRG flows should allow us to significantly further our understanding and quantitative access to the physics of strongly correlated systems. There are many applications that are natural extensions of that presented here. Here we only name a few of them that are already under investigation: (1) The reparametrisation potential of PIRG flows can be used to construct fully gauge invariant flows, and this possibility originates in the very fact that the composite field effective action is not identical to that in the original fields. This idea is discussed further in [62]. (2) Ground state expansions are highly interesting for an application to non-equilibrium evolutions of closed and open quantum systems, including systems with an initial far-from-equilibrium state. One may even base expansion schemes on time-dependent ground state expansions. (3) The flexibility of the choice of the target action allows for the systematic inclusion of results from other approaches such as (resummed) perturbation theory, hard thermal and dense loop results of low and high order, or density functional theory. (4) Finally, a combination of several of the present advances will be applied to the ongoing endeavour of resolving the physics of QCD at finite temperature and density. A specific first application is the phase structure of QCD, including the location of the critical end point with full apparent convergence and a small systematic error. We hope to report on these applications in the near future.
@@ Appendix A Symmetries and cumulants-preserving physics-informed RG flows
[…]
@@ A.1 Cumulants-preserving PIRGs
[…]
@@ A.1.1 Cumulants-preserving Wetterich flow
[…]
@@
Moreover, for $k\to\infty$, all theories tend towards the Gaußian theory with the momentum-dependent mass $R_{\phi}$, and the normalisation there is simply ${\cal N}_{W,k\to\infty}=1$. In combination this entails
@@ A.1.2 Cumulants-preserving PIRG flow
with a $\boldsymbol{g}$-independent constant on the right-hand side of 146. Finally we remark, that although the coupling dependence in 146 has to cancel, the specific value of the single terms $\dot{\mathcal{C}}_{k}(\boldsymbol{g})$ and $\Delta\dot{\mathcal{C}}_{k}(\boldsymbol{g})$ is not fixed for a specific set of coupling parameters $\boldsymbol{g}$. This freedom opens the gateway for optimisation procedures. The condition used in the present work is discussed in the following Section.
@@ A.4 Symmetry of the field transformation in the classical target action flow
Figure 9: Flowing field transformation $\partial_{t}\phi=\dot{\phi}$ for different choices of boundary conditions. One corresponds to the upper integration bound in 108 which is set to $\infty$ (or -$\infty$ for the dashed lines), whereas $c_{1}$ parametrises the constant part of the flow 107. It implements the choice of constant discussed in Appendix A. The blue curve indicates the value of $c_{1}$ for which $\dot{\phi}(\phi=\pm\infty)=0$ and 148 are fulfilled. The constant is a scale-dependent quantity; the figure uses $k=0.5$ as an example for the initial conditions specified in Section V.1.
@@ B.1 Reference solution
$Z_{\phi}^{(c)}(J_{\phi})=\int d\hat{\varphi}\,e^{-S[\hat{\varphi}]+J_{\phi}\hat{\phi}}\quad\textrm{with}\quad{\cal N}^{(c)}=1\,$, (153)
@@
$\int d\hat{\varphi}\,\hat{\varphi}^{2n}e^{-S[\hat{\varphi}]}$= $\,(-2)^{2n}\frac{d^{2n}\int d\hat{\varphi}\,e^{-S[\hat{\varphi}]}}{d\mu_{\varphi}^{2n}}$
@@
$\hskip-34.14322pt=-\frac{2}{\mu_{\varphi}}\langle\hat{\varphi}^{2}\rangle^{(c)}-\frac{8x}{\mu_{\varphi}}\frac{K_{\frac{5}{4}}\left(x\right)}{K_{\frac{1}{4}}\left(x\right)}\partial_{x}\log\frac{K_{\frac{5}{4}}\left(x\right)}{K_{\frac{1}{4}}\left(x\right)}\,$, (160)
@@ B.2 Benchmark with PIRG flows
In the present case we consider cumulants-preserving variants of classical target action, recall the discussion in Appendix A. In particular we implement 148, for which we have verified 146. The computation is initiated at $\Lambda=\exp(5)\approx 150$ and the RG-time is integrated up to t=15t=-15, corresponding to $k_{\mathrm{fin}}=4\times 10^{-5}$. The map $\phi(\varphi)$ is reconstructed using 55 and depicted in Figure 10(a). Using the map we can also depict the potential in terms of the original field $V_{\mathrm{cl}}(\phi(\varphi))$. This is shown in Figure 10(b), where we have explicitly omitted the constant shift in the target action 96 for better visibility. One can clearly discern 68 in the Figure: The solutions for the effective actions are genuinely different for different composites $\hat{\phi}$.
@@
which is an algebraic equation in $f_{0}$. We immediately find that for the classical target action flow, i.e. constant $\partial_{t}V_{T}$, $f_{0}$ has a non-trivial singularity where $2\rho V_{T}^{\prime}-\frac{2}{d+2}\frac{A_{d}k^{d+2}}{k^{2}+V_{T}^{\prime}}=0$. This singularity can be lifted by choosing $\partial_{t}V_{T}$ accordingly
@@ Appendix E Feed-down flows in the first order of the derivative expansion for O(N)-theories
$\hskip-11.38092pt+\frac{\phi^{a}\phi^{b}}{2\rho}\,\frac{1}{k^{2}+V^{\prime}_{W}(\rho)+2\rho V^{\prime\prime}_{W}(\rho)}\,.$ (182)